(A) write an equation for the line f that passes through (5,3) and has a slope of -8/9
(B) what is the slope of a line perpendicular to line f?

Found 2 solutions by amarjeeth123, macston:
Answer by amarjeeth123(575)

Let the line be y=mx+c where m is the slope and c is the y-intercept
y=-8/9x+c
It passes through (5,3)
Substituting we get,
3=-40/9+c
c=67/9
The equation of the line is 9y=-8x+67
8x+9y=67.
The slope of a line perpendicular to f is 9/8.

Answer by macston(5194)

PROBLEM A: EQUATION
y=mx+b slope=m, x=5, y=3, m=-8/9: solve for b
3=8/9(5)+b
3=(40/9)+b Subtract 40/9 from each side.
3-4.444=b
-1.444=b
We now have m=-8/9 and b=-1.444
ANSWER The equation for the line is f=-8/9(x)-1.444
PROBLEM B: A perpendicular line has slope that is negative reciprocal of the slope of the line. ANSWER A perpendicular line has a slope of 9/8.